# Machine Learning and the Physical Sciences

### v1

@@ 1 Introduction
In physics and engineering, the most impactful models are those that remain reliable even beyond the training data or observation domain, where controlled experiments or simulations are expensive, sparse, or beyond reach. Yet, the ability to master extrapolation remains out of reach for state-of-the-art machine learning models, including the latest forecasting approaches toner2025performance ; liu2024moiraimoe ; pucher2025evaluating ; li2025tsfm such as Foundation Models (FMs). FMs jin2024time ; woo2024unified are transformer-based neural networks with a considerably large number of parameters (hundreds of millions) pretrained on large datasets from diverse time series domains, but still have been shown to be outperformed by simple linear or seasonal models toner2025performance in extrapolation or long-range forecasting settings. More generally, deep learning models contrast with physical models having a symbolic structure, in the range of data regimes they can encode or generate. Indeed, in the context of fluid dynamics for instance, training a neural network solely on laminar flow data would lead to highly inaccurate predictions fukami2021model ; sutter2021robust ; chuang2020estimating for turbulent flows. Whereas a symbolic differential equation such as the Navier-Stokes equation- or its discretization in space, represents a very flexible data-generating process that covers a very wide range of regimes. This indicates a fundamental difference between symbolic models and neural networks and raises the question of characterizing this difference precisely to identify and design models that would extrapolate effectively. Although neural networks are typically over-parameterized and have a particular structure involving composing the same type of functions repeatedly, they are still approximators with explicit expressions, just like symbolic models, hence it is not clear what qualitative properties make them less suitable for extrapolation. Consequently, we tackle this question by making the following contributions:
@@ 3 Measuring Structural Variability
    The minimal order of the ODE: the degree of the minimal multivariate polynomial defining the ODE represents the most elementary discriminating quantity to consider, and encodes coarse qualitative change. For instance, oscillatory behavior cannot be observed in solutions of first order ODEs and require a second order term, as illustrated in figure 2.
@@
$f_{\theta}(x)=f_{1,\theta}(x)+\dots+f_{k,\theta}(x)$
Assume $(f_{j,\theta})_{j\leq k}$ do not span the variability classes, then for all $M>0$, there exists a smooth function $g_{M}$ such that
$\inf_{\theta\in\mathbb{R}^{p}}\|f_{\theta}-g\|_{\infty}>M$
@@ Appendix B Proof of proposition 1
Assume the variability deficit above holds. Then for every $M>0$ there exists a function $g_{M}\in C^{\infty}(K)$ such that
$\inf_{\theta\in\mathbb{R}^{p}}\ \|f_{\theta}-g_{M}\|_{\infty}\;>\;M.$
@@
$\inf_{\theta}\|g_{0}-f_{\theta}\|_{\infty}\;\geq\;\frac{|\Lambda(g_{0})|}{\|\Lambda\|}\;=:\;\varepsilon_{0}\;>\;0.$
Now fix $M>0$ and define $g_{M}:=\alpha g_{0}$ with $\alpha:=(M+1)/\varepsilon_{0}$. Then, for all $\theta$,
$\|g_{M}-f_{\theta}\|_{\infty}\;\geq\;\frac{|\Lambda(g_{M}-f_{\theta})|}{\|\Lambda\|}=\frac{|\Lambda(g_{M})|}{\|\Lambda\|}=\frac{\alpha|\Lambda(g_{0})|}{\|\Lambda\|}=\alpha\varepsilon_{0}=M+1>\;M.$
Hence $\inf_{\theta}\|f_{\theta}-g_{M}\|_{\infty}>M$, as claimed. ∎

### v2

@@ 1 Introduction
In physics and engineering, the most impactful models are those that remain reliable even beyond the training data or observation domain, where controlled experiments or simulations are expensive, sparse, or beyond reach. Yet, the ability to master extrapolation remains out of reach for state-of-the-art machine learning models, including the latest forecasting approaches toner2025performance ; liu2024moiraimoe ; pucher2025evaluating ; li2025tsfm such as Foundation Models (FMs). FMs jin2024time ; woo2024unified are transformer-based neural networks with a considerably large number of parameters (hundreds of millions) pretrained on large datasets from diverse time series domains, but still have been shown to be outperformed by simple linear or seasonal models toner2025performance in extrapolation or long-range forecasting settings. More generally, deep learning models contrast with physical models having a symbolic structure, in the range of data regimes they can encode or generate. Indeed, in the context of fluid dynamics for instance, training a neural network solely on laminar flow data would lead to highly inaccurate predictions fukami2021model ; sutter2021robust ; chuang2020estimating for turbulent flows. Whereas a symbolic differential equation such as the Navier-Stokes equation- or its discretization in space, represents a very flexible data-generating process that covers a very wide range of regimes. Furthermore, symbolic learning has demonstrated stronger extrapolation properties in biological applications dakhmouche2024robust ; dakhmoucherobust . This indicates a fundamental difference between symbolic models and neural networks and raises the question of characterizing this difference precisely to identify and design models that would extrapolate effectively. Although neural networks are typically over-parameterized and have a particular structure involving composing the same type of functions repeatedly, they are still approximators with explicit expressions, just like symbolic models, hence it is not clear what qualitative properties make them less suitable for extrapolation. Consequently, we tackle this question by making the following contributions:
@@ 3 Measuring Structural Variability
    The order of the ODE: the highest order derivative of the ODE represents the most elementary discriminating quantity to consider, and encodes coarse qualitative change. For instance, oscillatory behavior cannot be observed in solutions of first order ODEs and require a second order term, as illustrated in figure 2.
@@
$f_{\theta}(x)=a_{1}f_{1,\theta}(x)+\dots+a_{k}f_{k,\theta}(x)$
Assume $(f_{j,\theta})_{j\leq k}$ do not span the variability classes, then for all $M,\varepsilon>0$, there exists a smooth function $g_{M}$ such that
$\sum_{i=1}^{n}\|Y_{i}-g_{M}(X_{i})\|_{2}\leq\varepsilon\quad\textrm{and}\quad\inf_{\theta\in\mathbb{R}^{p}}\|f_{\theta}-g_{M}\|_{\infty}>M$
@@ Appendix B Proof of proposition 1
Assume the variability deficit above holds. Then for every $M,\varepsilon>0$, there exists a smooth function $g_{M}$ such that
$\sum_{i=1}^{n}\|Y_{i}-g_{M}(X_{i})\|_{2}\leq\varepsilon\quad\textrm{and}\quad\inf_{\theta\in\mathbb{R}^{p}}\|f_{\theta}-g_{M}\|_{\infty}>M$
@@
$\inf_{\theta}\|g_{0}-f_{\theta}\|_{\infty}\;\geq\;\frac{|\Lambda(g_{0})|}{\|\Lambda\|}\;=M+2>\;0.$
Data correction. Let $\{x_{i}\}_{i=1}^{n}\subset K$ be the sample inputs. Choose pairwise disjoint open neighborhoods $U_{i}\Subset K$ and $\phi_{i}\in C_{c}^{\infty}(U_{i})$ with $\phi_{i}(x_{i})=1$ and $\|\phi_{i}\|_{\infty}\leq 1$. Define coefficients
$t_{i}:=\operatorname{clip}\big(y_{i}-g_{0}(x_{i}),\,\varepsilon/n\big),\quad\text{i.e.}\quad t_{i}=\begin{cases}\phantom{-}\varepsilon/n&\text{if }y_{i}-g_{0}(x_{i})>\varepsilon/n,\\ -\varepsilon/n&\text{if }y_{i}-g_{0}(x_{i})<-\varepsilon/n,\\ y_{i}-g_{0}(x_{i})&\text{otherwise,}\end{cases}$
and set the correction
$r(x):=\sum_{i=1}^{n}t_{i}\,\phi_{i}(x),\qquad g:=g_{0}+r.$
Because the supports are disjoint and $\|\phi_{i}\|_{\infty}\leq 1$,
$\|r\|_{\infty}\;=\;\max_{i}|t_{i}|\;\leq\;\frac{\varepsilon}{n}\;<\;1.$
At the data points we have $g(x_{i})=g_{0}(x_{i})+t_{i}$, hence
$|y_{i}-g(x_{i})|=|y_{i}-g_{0}(x_{i})-t_{i}|\leq\frac{\varepsilon}{n},\quad\Rightarrow\quad\sum_{i=1}^{n}|y_{i}-g_{0}(x_{i})|\leq\varepsilon.$
Margin to the model class. For any $s\in S:=\ker(T)$, by the triangle inequality,
$\|g-s\|_{\infty}\;\geq\;\|g_{0}-s\|_{\infty}-\|r\|_{\infty}.$
Taking the infimum over $s\in S$ and $\|r\|_{\infty}\leq 1$,
$\mathrm{dist}(g,S)\;\geq\;\mathrm{dist}(g_{0},S)-\|r\|_{\infty}\;>\;(M+2)-1\;>\;M,$
Since $\mathcal{F}\subset S$, it follows that
$\inf_{\theta}\|g-f_{\theta}\|_{\infty}\;\geq\;\mathrm{dist}(g,S)\;>\;M,$
which proves the claim. ∎